At winter's end, Lake Superior's 82,000-km2 surface is frozen to a depth of 1.3 m. The density of ice is 917 kg/m3.

How much energy does it take to melt the ice?

If the ice disappears in 3 weeks, what is the average power supplied to melt it?
Answer #1

here,

Volume of ice = 82000 * 1.3

Volume of ice = 106600 m^3

mass of ice = volume * density

Mass of ice =  106600* 917

Mass of ice = 977522000 Kg

then

Enegy needed to melt ice = 977522000 * 334 KJ

Enegy needed to melt ice =3.265 * 1010 KJ

energy it take to melt the ice is 3.265 * 1010 KJ

Power = energy /time

Power =  3.265 * 1010 / (3* 7 * 24 * 3600)

Power = 17994.5 kW